A generic iron preparation has a concentration of 20 mg/ml. Calculate the number of milliliters required to administer 200 mg/day for 180 days.
Answer #1

ANSWER.

* Calculation of number of ml of drug administered for 180 days.

       Formula

           Volume(ml) = Desired dose /Dose in hand * Quantity.

      Here

             Desired dose (D) = 200 mg/day.

             Dose in hand (H) = 20 mg.

                  Quantity (Q) = 1 ml.

   As per above formula

                 Volume(ml) = D/H * Q

                                    = 200/20 * 1

                                    = 10 * 1

                                    = 10 ml/day.

     Volume(ml) of drug required for 180 days = 180 * 10

                                                                        = 1800 ml /180 days.